1. Solve the following simultaneous equations by using Cramer’s rule

3x+2y=13
2x-y=4
1
Expert's answer
2013-01-16T11:10:13-0500
We need to find 3 determinants for matrices

| 3 2| |13 2 | |3 13|

| 2 -1| | 4 -1 | |2 4|


for the first D=3*(-1)-2*2=-7

for the second Dx=13*(-1)-2*4=-21

for the third matrix Dy=3*4-2*13=-14

Then by Cramer's rule x=Dx/D=3, y=Dy/D=2 so x=3,y=2

If you have problems with understanding Cramer's rule please check the video below. We shot it to discuss and explain Cramer's rule application.